OpenStudy (anonymous):

How many pounds of walnuts that cost $7.60 per pound must be mixed with 6 pounds of cashews that costs $12 per pound to make a mixture of nuts that costs $9.25 per pound? Select one: a. 10 pounds b. 12 pounds c. 10.5 pounds d. 8 pounds

OpenStudy (yanasidlinskiy):

The final cost is = Total money / Total mass of nuts = 9.25 The money for the cashews is, 6Pounds X (12$ / Pound) = 72$ The money for the Walnuts is, xPounds X(7.6$ / Pound) = 7.6x$ Total Money is: 7.6x$ + 72$ Total mass is: Mass of cashews + Mass of Walnuts = 6 + x Total Money / Total mass = 9.25 = (7.6x$ + 72$) / (6 + x) Solving the equation: 9.25(6+x) = 7.6x + 72 Note: "$" is not necesary Then 55.5 + 9.25x = 7.6x +72 9.25x -7.6x = 72 - 55.5 1.65x = 16.5 then x= 10 Pounds

OpenStudy (anonymous):

thnx(:

OpenStudy (yanasidlinskiy):

Your Welcome:DDDD